OpenStudy (anonymous):

How to find x in his solution: 4sinxcosx=1

Parth (parthkohli):

Divide both sides by \(2\).\[2\sin(x)\cos(x) = \dfrac{1}{2}\tag{*}\]Can you rewrite the left-hand-side of \((*)\)?